If A+B=45 the [cotA-1 ][cotB-1] is 
a).1 b). ½ c). -1 d). -2 d). 2

Grade:11

1 Answers

Utsav Basu
6 years ago
A + B = 45
So, 
cot 45 = cot (A + B) = (cot A * cot B – 1)/(cot A + cot B)
or, cot A + cot B = cot A * cot B – 1
or, cot A * cot B – (cot A + cot B) = 1
Now, (cot A – 1)(cot B – 1) = cot A * cot B  – (cot A + cot B) + 1 = 1 +1 = 2